40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog, Page 37 of 61  Not logged in ELOG logo
ID Date Author Type Category Subject
  1249   Thu Apr 30 17:53:20 2015 ZachCryostatPurchasesGetters gotten

Being the go-getters we are, we got our spare getters gotten. They'll be stored in the cabinet for when we need them.

 

  1248   Fri Apr 24 04:31:48 2015 ZachLaserSiFiCavities locked simultaneously, transmission setup built

I was running into some issues optimizing the servos with the present actuation setup, and I had a conversation with Dmass about another short-term solution (more on that at the end), so I figured I'd try a simpler approach to getting both beams at the output for the time being: I knew I could get one cavity locked with a simple SR560 servo, so I decided to try locking the other in the same way. I didn't expect this to work very well, since the other diode driver I had to work with only has ~1-kHz bandwidth, but miraculously it did. The lock is a bit shaky, but it holds indefinitely and relocks readily as with the other cavity. I guess there's just enough less noise with the cavities inside the cryostat.

With the cavities locked, I went about building the transmission setup. The beams are both directed west via 90R/10T mirrors, where they then each pass through a QWP to return them to linear polarization and then a HWP to set them to S, which is the favorable polarization for optimal 50/50 splitting with the beamsplitter we are using. The waveplates were optimized using an analyzer PBS, and I got the polarization contrast down to about 0.3% in both cases. I used geometry to make sure the beams were combined at the same gouy phase. One ouptut of the beam-combining BS is then steered and focused onto a 1611 PD, and the other output will be put on an auxiliary PD for DC TRANS diagnostics. I also decided to overload the CCD camera as a dual-beam sensor by just pointing both of the leakage beams through the first steering mirrors at the camera, side by side (as seen in the photo above). The beams are positioned relative to each other on the screen as the cavities are seen from the rack (E beam on the left, W beam on the right).

The whole setup looks a bit like Stiltsville right now (explained here for non-Floridians), which isn't the long-term plan. When I get a feel for how much space I really want, I'll get a breadboard and just elevate it by the ~2" required. An added benefit to that scheme is that the whole thing can be removed in one piece for access to the cryostat.

Here's a glamour shot of the I/O optics:

I aligned the combined beams and tried to get my first transmission beat, but didn't have any luck right away and needed to call it quits. The procedure I had planned for doing this in the future is:

  1. Use the frontal beat to bring the two beams close together in frequency. One easy place to do so is at the RIO test report temperatures.
  2. Scan one beam until a resonance is found and then lock.
  3. If the frontal beat has been lost (fairly likely, as the frontal beat PD is an 1811 with BW ~ 125 MHz, while the cavities have an FSR of >2 GHz), scan the second laser until the beat is found.
  4. Once beating, scan the second (still unlocked) laser in both directions and find the closest resonance, then lock.
  5. In the worst case, the two cavities' closest modes are ~1 GHz apart (i.e., FSR/2), so it should be possible to see a beat if the procedure is done properly. For the real experiment, we'll have some slow length actuation (e.g., via temperature) so that we can bring the two cavity modes as close together as we want.

Returning to the actuation situation, Dmass actually prefers that I hijack his diode drivers temporarily, rather than trade for his ITC510 unit (even though he is not using the diode driver side of that unit, he is worried about having to retune the temperature loops with a new unit). This is actually way better for me, as well, since it will really let me tune up my servos the way I'll want them when I grow up and get my own Rich drivers. Now that the optical layout is more or less complete (at least for the short term), I can focus almost exclusively on the loops. There are some spare long BNC->DSub cables for exactly this purpose that I can use to drive my lasers without moving the drivers from the CryoCav rack, so I'll get on that shortly.

 

  1247   Thu Apr 23 03:09:49 2015 ZachDailyProgressSiFiWindows installed, cryostat top with cavities placed in body, cavities realigned, locked

These things happened over the last few days but haven't been elogged.

Windows installed

I installed the AccuGlass windows into the flanges, but ran into some trouble as one of them cracked. It happened on the 4th one, and I did not use any more force than seemed reasonable given that the glass-on-o-ring contact has to hold the vacuum. I think we might need to get some teflon gaskets to soften the metal-glass contact.

 

Cavities installed into chamber body, realigned, locked

Despite the above, since we had 5 windows, I finished putting them on the cryostat and went about installing the cavities (mounted to the coldplate and cryostat top) into the main body. I transferred the top from the open-air test rig (see CRYO:1244) into the chamber body, then positioned the whole assembly according to some reference marks I made on the table. Using a fast lens in transmission of one of the cavities, I was able to align the cryostat well enough by hand to see some high-order flashing (the test rig is ever-so-slightly taller than the assembled chamber, so what I saw was mostly TEM0n modes). With some tweaking of the input alignment, I aligned each beam to its cavity's TEM00. Here is one cavity locked while in the chamber:

It was all very deterministic, so that inspires some confidence in the procedure.

Other stuff

I also tuned up the W RFPD to match the temporary testing TF I set on the E one (readout 30 MHz, notch 60 MHz---see CRYO:1242), and used it to lock one of the cavities.

I attempted to modify the PDH2 from gyro configuration into a useful one for testing here, but I may have changed too many variables at once; I knew I wanted some more gain in the ~1 kHz region, so I used an extra P/Z than what is available with the uPDH box to do so. Somehow, I'm not getting any good locking action even though the TFs are identical at DC and near the UGF of the known working loop at ~30 kHz.

Also, I remembered that there's no hope to do any good locking with the other (standalone, LDC201C) laser driver that I have, since its bandwidth is about a kHz (CRYO:1205). I'm asking Dmass if it's OK if I trade him my standalone temperature controller (TED200C) for one of the integrated units (like the one I have one of), since he's only using the temperature controller half.

  1246   Tue Apr 21 21:30:50 2015 ranaLab InfrastructureDrawingsLab Layout circa Feb. 2015 (according to Nic)

Heavy Duty (14 gauge steel) cabinet w/o glass doors, but with door storage ordered.

From CryoLAB: April 21, 2015
  1245   Tue Apr 21 14:19:16 2015 ZachLaserPurchasesFiber isolators arrived

We bought some IO-H-1550APC fiber isolators from ThorLabs last week. They arrived today:

  1244   Fri Apr 17 05:15:26 2015 ZachLaserSiFiTest cavities rebuilt on cold plate, aligned, locked (separately)

Given the level of noise I saw in air with the table-mounted test cavity yesterday (CRYO:1242), I decided to move straight into in-vacuum testing.

The first thing I needed to do was set up an in-air support system for the top half of the cryostat for mounting and alignment (i.e., I needed to suspend the cold plate in the position it will be in with the cryostat closed, but with access to the work area). By what I can only describe as a magical coincidence, the height of the upper/lower joint on the cryostat is almost exactly the height of two long 1" pedestals stacked together, and there are 8-32 taps for the cryostat joining that can be used to mount them (huzzah!):

I temporarily stored the getter in a ziploc bag for the duration of this preliminary testing, then positioned the lifted coldplate over the test cavity area. Then, I went about building the cavity in sleeping-bat position on the cold plate. The cryostat windows are 149 mm above the table (145 mm from the bottom of the cryostat, but lifted another 4 mm by the caps of the screws holding the floor on, which are not countersunk). Before building the cavities, I assembled perisocpes to lift the beams from 4" to 149 mm using discrete components.

 

I was a bit too optimistic about space when building the test cavities on the table, so I needed to shorten them a bit to around 2.5". Because of this, and because the distance to the waists changed, I needed to recalculate the MMTs. The solution only shifted the lenses by a few inches:

------------------------------------------------------------
Other solution:

mismatch: 0.00015463
w0x = 200.5644 um 
w0y = 200.5644 um 

lens 1: f = 103.2118 mm
lens 2: f = 103.2118 mm
Distances:
d1 = 15.2397 cm
d2 = 25.9767 cm
d3 = 74.0136 cm
(Total distance = 115.23 cm)

After realigning the E cavity, it locked again with no trouble using the same servo and settings (note: this shot was actually taken before building the W cavity and periscope):

The lock was actually quite a bit stabler seeming than yesterday's, perhaps due to some mechanical low-passing from the support system. I'm hoping this gets better, not worse, with the full cryostat in place.

After building the W cavity and periscope and aligning, I borrowed all optoelectronics from the E path (criss-crossed laser fibers, swapped the E RFPD in, and used the same electronics chain) and it, too, locked with little trouble, though it seems like the modematching is somehow not as good on that path yet. Here are preliminary control spectra from each cavity (note that these were taken at different times, which I'm hoping accounts for some of the slight differences in noise features):

The control spectra are plotted alongside the michelson displacement noise measurement from the other day (CRYO:1238) as well as the PMC-derived laser frequency noise measurement from February (CRYO:1205). Some things to note:

  • Admittedly, the W cavity measurement was taken hastily at the end of the night. It looks like the UGF was rather low and some gain peaking accompanied it at a few kHz.
  • By and large, the displacement noise looks better with these cryostat-mounted cavities than with the table-mounted michelson. In particular, there seems to be some filtration above a few hundred Hz, which I think greatly improved the locking stability today vs. yesterday on the table.
  • The E cavity spectrum, which is more trustworthy and which should have been made with a UGF near 100 kHz, intercepts the expected laser frequency noise level above a kHz or so.
  • Note that the red trace sits on the dark noise level above 4 kHz, and its being near the laser frequency noise level on this plot is pure coincidence (the michelson wouldn't have been sensitive to this).

All in all, this looks pretty reasonable and things are promising for the in-vacuum testing. yeslaugh

Here is a parting photo of the new layout:

For tomorrow:

  • Install cryostat windows
  • Close cryostat and pump
  • Tune W RFPD and PDH2
  • In-vacuum simultaneous lock
  • Build TRANS setup with beat
  1243   Fri Apr 17 04:39:25 2015 ZachUpdateSiFiTwo 150mm wafers sent for dicing

I sent two of our 150mm (~6") wafers to be diced by American Precision Dicing (Justin's recommendation). One wafer will be cut into 100mm x 10mm rectangles, and the other into 50mm x 10mm rectangles.

Justin gave me a nice wafer holder he didn't need (left), so I've stored all but the two I shipped in that one, and shipped the two wafers using the crappier one they came in from University Wafer (right):

While I had the holders open, I took this shot of the reflection from one of the wafers to demonstrate the niceness of the polish:

  1242   Thu Apr 16 15:58:02 2015 ZachLaserSiFiGyro RFPDs installed, tuned, E cavity PDH tuned and cavity locked

Well, I got partway through the plan for the day, anyway.

Rather than align the W cavity as I had done the E, I decided to continue working on locking the E cavity so that I could copy all the work wholesale to the other side at the end. The first thing I did was to install the gyro RFPDs. To facilitate this, I installed a NIM crate in the bottom of our rack (one of the spares that was stored under the gyro table). This will be used to power the PDs, the PDH2 board, and some of my homebrew filters/preamps if necessary. After powering what became the E REFL PD up, I installed it and focused the REFL beam onto it.

 

I checked to see if either of our crystal oscillator frequencies were within the tuning range of the PD as it's stuffed now. They weren't, so I just tuned it to 30 MHz (and notched 60 MHz) for temporary testing. Recall that these are aLIGO-style PDs that don't have maximal gain at the resonant readout frequency. In practice, they are tuned by looking at the notch seen by the anode, rather than looking at the readout node as seen below. (Note: the delay seen is consistent with the optical path lengths through the fiber and free space).

I then did some playing around with an SR560 servo, moving the pole and gain until I got some weak locking action. Then, I systematically searched the parameter space until I found the best stability (still bad, though). Using that information, I built into the uPDH box (#1437) a TF that had similar gain in the UGF target region of ~50-100 kHz, but much more low-frequency gain. This ended up being something like zpk([10k,10k], [50, 50], 1000), where the cavity pole at ~40 kHz returns the loop to 1/f above there. I would have put one of the servo zeros at this frequency, but I had too large a capacitor for it to make sense due to the smallness of the resistor needed---this is something I can change if necessary.

Plugging that bad boy in and playing with some attenuation before and after the servo, I got a reasonably stable lock, but nothing too stellar. It holds for many minutes, but it is in a delicate balance (often not very balanced) between lots of unsuppressed audio noise and some instability in the ~100-kHz+ band due to the plant. This is exacerbated by the low bandwidth of the driver (see CRYO:1205---I'm using the ITC502 at the moment). I tried doing some feedback using the bias tee input, but I wasn't able to lock at all with this method. Maybe I need to do some high-frequency crossover to it while keeping the low-frequency actuation going through the driver. Here is a shot showing the transmitted beam on the camera and card, as well as a scope trace of a lock acquisition:

 

In the above, GREEN is TRANS, CYAN is error, and MAGENTA is actuation. The error signal is dominated by periodic oscillations at >100 kHz, while the transmission shows plenty of audio-band noise.

Given the measurement that I made the other day on essentially the same physical system (CRYO:1238), it looks like the audio-band RMS is on the order of almost a nanometer, which is ~50x the linewidth of this cavity. So, I need more gain or less noise at 1 kHz. There are two options:

  1. Switch to the PDH2 and add rapid rollup below ~30 kHz with up to 4 P/Zs
  2. Just go to vacuum already, since we don't really need to contend with this much noise in the end

Since I want both sides to be roughly balanced (i.e., same loops, etc.), I'm tempted to just do (2). I've already got everything pretty well aligned and I know the electronics TFs are not crazy, so I think that's what I'll do.

Quote:

For tomorrow:

  • Repeat the last step above for the W cavity
  • Install and tune the gyro RFPDs
  • Determine appropriate PDH servo TFs and modify boards
  • Lock cavities
  • Build transmission beat setup

 

  1241   Wed Apr 15 01:19:14 2015 ZachLaserSiFiInput paths, frontal beat, test cavities set up and aligned, ready to lock

I bought and received the last optics I was missing (90R/10T BSs) last week, so I began building the real experiment today. Here is what it looks like so far:

Frontal beat

The first thing I did was build the frontal beat, which we will use to locate and adjust the beams' frequency offset without having the cavities locked. This is fit in a rather compact layout between the main beams, using a 10% pickoff from the second mirror in each path. The beat alignment DOFs are matched by the actuation offered by 1) the extra steering mirror in the E (left above) beam path and 2) the combining beamsplitter. A further steering mirror puts the beam on the PD after a focusing lens, and the (small) reflection is dumped. Here are the beams beating near 50 MHz:

Test cavities

We want to use some dummy cavities to set up and test our electronics. To make these, we've used some of Dmass's spare fused silica ATF-coated optics (coating run V6-593/594---scans attached). I chose to use two 50-cm mirrors for each cavity, with a length of ~4" (this is just roughly as big as you can make it with two big mirrors within the vacuum area---note the cryostat boundaries drawn onto the table). The transmission of these mirrors is 0.016%, giving a finesse of ~20k. Here are the arbcav plots:

 

I temporarily relocated the cryostat to the central table so I could put the dummy cavities within its boundaries on the main table. Installing only the cavity end mirrors first, I finished the main beam paths, aligning the beams along the holes and level at 4" for the main stretch, then installing the steering mirror zigzags to bring the beams in to the cavity longitudes. With the beams centered on the cavity output mirrors, I then calculated a modematching solution. Modematchr came up with plenty of options, but I chose this one because it used only f=100mm lenses (which I had bought specifically for mode matching) and left plenty of room near the cavities for the circulating optics:

------------------------------------------------------------
Other solution:

mismatch: 0.0002397
w0x = 222.7609 um 
w0y = 222.7609 um 

lens 1: f = 103.2118 mm
lens 2: f = 103.2118 mm
Distances:
d1 = 11.8781 cm
d2 = 26.4822 cm
d3 = 72.1297 cm
(Total distance = 110.49 cm)

As you can see, the predicted best mismatch is 0.02%.

I installed these (on slotted bases, so they can be adjusted), then verified that the output beams were roughly as expected---they were. I then used the cavity output mirrors to retro-reflect the beams, which served as a further modematching sanity check (since the retroreflected beams agreed transverse-spatially). Next, I mounted and installed the circulating optics (PBSs and QWPs), tuning the initial HWP and the QWP to maximize forward transmission and backward rejection, respectively. Finally, I focused the E REFL beam on a PDA255 for temporary testing.

Now the input beams and end mirrors were aligned, so all I had to do was install and align the cavity input mirrors. Before doing so, I borrowed the 1550nm-sensitive CCD camera from Dmass's setup and placed it behind the E cavity end mirror. With all the lights off, I could very faintly see the transmitted beam, and I centered the camera onto it. The REFL PD gave me a good reference, so just installed the input mirror and directed the prompt reflection back to the PD. Immediately upon doing this, I saw strong transmission flashes on the camera.

Tweaking the cavity and input beam alignment somewhat while scanning the laser frequency, I did a rough TEM00 maximization. The REFL dips indicate ~80% coupling, which I think is as good as I'll bother going before locking. yeslaugh

 

For tomorrow:

  • Repeat the last step above for the W cavity
  • Install and tune the gyro RFPDs
  • Determine appropriate PDH servo TFs and modify boards
  • Lock cavities
  • Build transmission beat setup
Attachment 7: V6-593_Ab450.pdf
V6-593_Ab450.pdf
Attachment 8: V6-594_AB450.pdf
V6-594_AB450.pdf
  1240   Tue Apr 14 22:27:18 2015 ranaLab InfrastructureDrawingsLab Layout circa Feb. 2015 (according to Nic)

Some cabinet options (for storage of heav-ish things):

  1. SC-4824 (48" W, 24" deep, 78" H) from TecLab
  2. L47HB from Hanson Lab (same as above, but with glass doors and nicer hinges and rubber door stoppers)
  3. 73324 from hemcocorp.com, same but only 18" deep.
  4. A few on Amazon for ~1200$: http://amzn.com/w/1J0TVPU10HHSN
Attachment 1: CryoLabLayout_150303.png
CryoLabLayout_150303.png
  1239   Mon Apr 13 11:24:32 2015 ranaNoise HuntingMechanicsPost and base vs. pedestal and fork

Its a nice result. As you say, we are lacking in good writeups about this topic. Mostly they're plots in the iLIGO elogs which were never collated and so are lost to the mists of time...

But I think this is missing one of the major points: the objection to the pedestal/fork combo has to do with the care required to assemble it. When assembled carefully, as you have done, it works well. But taking an ensemble of them in a larger lab has shown from observation, that some fraction of them are assembled poorly. Usually its the last step of nudging the fork which is not done or the final tightening pulls the fork back by a mm or so.

And I think all of these kinds of measurements are including a significant bias. To make a good comparison, we'd have to check that the torque used on the 1/4-20 screws is "good". Also, the finding from assembling the ISC tables back in the late 90's was that the torque used on the optic set screw is important. Too tight makes some distortion of the glass, too loose and you get some springiness. Another fork parameter which I believe is important is the distance between the pedestal and the screw. This removes some of the parameter spacei in the 'arbitrary' positioning capability of the pedestal/fork. When Mike Landry made these measurements ~2000, I believe he found that the springiness of the fork screw was a parameter in the extreme position cases.

Also, in the several previous incarnations of this experiment, we used a little computer speaker as a white noise source to drive the mount under test to make sure that the source noise was not changing.

  1238   Sat Apr 11 23:42:03 2015 ZachNotesMechanicsPost and base vs. pedestal and fork

Tonight, I got to do an experiment that I've wanted to do for some time now.

For years, I've heard in conversations with people who shall remain nameless (unless they care to contest this work) that the 3/4"-post-on-rectangular-1/4"-thick-base optic support method (used at the 40m and adopted into LIGO) is better than the standard 1"-pedestal-and-fork method (used by many experiments in our own labs and elsewhere). After many attempts, I have never succeeded in getting any hard data to support that claim. So, I decided to make a measurement myself.

I set up a simple michelson using one of the SiFi beams, once using each support scheme for the beamsplitter and end mirrors:

 

There is a HWP to find the polarization for which the "50:50" BS is closest to balanced, a lens to focus into the IFO, a second lens to focus the AS beam onto the PD, and the PD itself, which is a PDA255. There is an ND=0.7 filter on the PD.

Below is a scope screenshot of some fringing action when pushing on an end mirror. The contrast defect in each case was pretty low at ~2.5 x 10-3.

Once the IFO was aligned in each case, I pushed on one mirror a bit to creep it into a half-fringe state. This took some time, since I had to push and wait a few seconds for it to settle. After doing that, I took a spectrum (actually 3 at different spans from 1 kHz to 100 kHz). The results are below, with a zoomed plot to the right.

 

Conclusion

As you can see, the difference is pretty minimal. The post-and-base setup has slightly higher RMS below ~1 kHz, owing to two high-Q resonances (at 340 Hz and, to a lesser extent, 920 Hz). My detractors will accuse me of bias (e.g., in tightening, etc.), but I invite anyone to come test this with their post-and-base clamping chops.

With this, I'd like to put to rest the notion that the post-and-base method is somehow fundamentally superior. I DO acknowledge that there are definitely wrong ways to use the pedestal-and-fork, and this can lead to the non-idealities noted in the folklore. The post-and-base method is foolproof in a way, since the proper procedure is somewhat manifest (use two screws, use washers, etc.), while the pedestal-and-fork requires some diligence to get just right. However, with just a little bit of care up front, the pedestal-and-fork offers huge advantages:

  • Arbitrary placement (most importantly, the ability to always work along table hole axes, which makes alignment incredibly easier).
  • Free-hand alignment, since the fork can be placed without jostling of the optic (see below).
  • Space, since it takes up way less of it.

 

How to fork

To mitigate the potential recklessness of this post, I offer the Zach-Approved™ Forking Method.

First of all, this is the only fork you should be using, the Newport PS-F (maybe there actually are other acceptable ones, but none that I've found that don't apply horizontal forces on the pedestal upon clamping):

Now, the forking method (accompanied by the GIF below):

  1. Locate where you want to place the optic.
  2. Place the optic and align by hand. The pedestal is heavy enough to hold itself in place with friction.
  3. Locate the appropriate screwhole and set the fork gently onto the pedestal, slightly un-engaged. Of crucial importance is that you pick a hole that will be somewhere close to the middle of the slot. Working along the table hole axes ensures there should be several options in every case.
  4. Using your finger, engage the fork on the pedestal. If you use just enough force to move the fork, it will come to a stop when fully engaged and you won't have moved the pedestal (and optic) at all.
  5. Insert the screw and hand-tighten a few threadlengths.
  6. Use a ball driver to finish screwing almost until tight.
  7. Just before tightening, give the fork one final nudge against the pedestal. Sometimes I use the ball driver tip, but this can also be done with your hand so that you can keep the driver in the cap.
  8. Tighten as desired.

At no point after initial alignment should you have to touch the optic, and, if you follow the procedure above, the optic orientation should not have shifted by more than a mrad or so. You can see how little the optic moves over the operation in the GIF.

  1237   Sat Apr 11 19:03:58 2015 ZachElectronicsSensorsRL readout PD

Seeing Hartmut's talk at the last LVC meeting about innovative DC photodetector designs (something necessary for future squeezed IFOs) reminded me of some investigation I did into the same while at LLO. One thing I did a fair bit of work on while there was the DC current subtraction idea (c.f. LLO:6449 and 6532), but another thing I spent time modeling was the concept of using an RL network, as Hartmut is exploring now.

The circuit I was considering differs somewhat from Hartmut's idea. In his circuit (at left below), the inductor ("L1") and input resistor ("R1") perform a current branching: at low frequencies, the photocurrent sees low impedance to ground through the inductor, and therefore does not pass through the transimpedance amplifier and get converted into an output voltage; at high frequencies, the inductor looks like an open circuit, and all the current passes through the TIA. Ideally, this leads to an effective frontend whitening that allows for a high Z at audio frequencies. In practice, one would use either the DC resistance (DCR) of the inductor, or perhaps an extra resistor in series, to set the DC Z, which would be Z(0) = R_2 \frac{R_{DC}}{R_{DC} + R_1}, where RDC is the DC resistance of the inductor path. One problem with this design is that, since RDC cannot be arbitrarily low due to the DCR of the inductor, one must choose an R1 that is high enough to set the DC Z to a low enough level. Roughly speaking, this means that the value of R1 must be approximately the ratio of the desired AC and DC transimpedances (typically a factor of 100 or so), times RDC. Since RDC will be on the order of 100 Ohms, R1 must be on the order of 10 kOhm. This in turn means that the current noise of the amplifier is fully converted by this high impedance at all frequencies, which ruins the SNR of the detector at low frequencies (you could use a low-current-noise part, but then the voltage noise kills you directly).

The circuit I had in mind is at right below. As you can see, the amplifier in this case is only used as a unity-gain buffer for the passive readout circuit (though one could consider adding a switchable flat gain for low-current operation, as in the ZSWITCH feature of the currently used DCPDs). This design works simply by having a passively different transimpedance at different frequencies: at low frequencies, the inductor shorts the large resistor and the transimpedance is just the inductor DCR plus the additional series resistance to ground (50 Ohms in the schematic); at higher frequencies, the impedance increases until it is limited by the parallel resistance (10 kOhms here, plus the series 50 Ohms). With this topology, the current noise always sees the same impedance to ground as the photocurrent does (i.e., the transimpedance), and there is no extra reduction in SNR. The "DC" section is not necessary in principle, and in fact it always has worse SNR for a finite inductor DCR, but it could be used as a calibration path for the DC response due to potential nonlinearity of the inductor.

 

As a side project, I've started doing some testing of this design. To start, I bought a ginormous 4-H inductor from DigiKey:

Transfer function

The first thing I did was to verify the transfer function. To do this, I biased one of our 3-mm diodes with the M2 circuit bias supply, then sent the anode into a breadboard version of the RL circuit. The parameters were slightly different: L = 4 H (DCR ~ 60 Ohm), RAC = 10k, RDC = 39 Ohm. I then put one of our SiFi lasers on it and modulated the power using its fiber amplitude modulator. Here is the result:

As you can see, it performs just about as expected. A couple notes:

  • The LISO trace above has been adjusted using the inductance and diode capacitance as fit parameters, since they are not known precisely a priori.
  • The slow upturn at low frequencies is fairly well explained by the amplitude modulator response at low frequencies (see CRYO:1187).

Noise

Of course, the biggest concern with using such a big inductor is the additional noise it might inject, particularly due to pickup. Below is a summary plot of some measurements I made on this circuit, together with some theory curves and the currently used DCPD for comparison.

Notes:

  • Obviously, there is a strong presence of pickup here. These traces are also the result of wrapping the inductor in metal and orienting it to minimize noise. Yes, the pickup is prohibitively bad as it appears here, but I have an idea to get rid of it (see below).
  • Below a few Hz and above ~1 kHz, the circuits behave as they should, except that the LT1128 exhibits excess current noise in the high-frequency region. This is no surprise, as we have never acheived the "typical" noise performance advertised by Linear (see ATF:1890---Rana points out that this could potentially be due to our not following proper electrostatic discharge practices, but it seems that they are always out of "typ" spec---not "max"---in the exact same way). In this case, though, the noise is even worse than the "max" level, which is an extra party foul. My proof of this is that the OP27-stuffed circuit does what it says it should at these frequencies. Note also that I had only some modest thermal shielding, and that's why the noise shoots up below a few 100 mHz.
  • Given the typical LT1128 performance in the real world, it may be that the OP27 is the ideal buffer amp for this circuit, since it adds a factor of a few at high frequencies (where we are most likely to be sensing noise limited). It's also only a factor of a few worse than the fantasy LT1128 curve at low frequencies.
  • For all its nascent imperfections, this is still already a detector with 10x better noise both at 1 Hz and above 1 kHz than what we have in aLIGO, so that's already promising. yes

Reducing the pickup:

So, we are left with the problem of being highly sensitive to a signal injected into the coil, but not to one induced by external fields. Luckily, the guitar industy has had a solution for almost exactly the same issue for about 80 years now: the humbucker. Of course, it's slightly different, since in the guitar case you want to be sensitive to an induced signal (i.e., the signal from the string, which is deliberately made differential-mode by reversing the polarity of the magnets inside the coils), but I believe the same principle should apply. In our case, we'll put the inductors in series electrically, but adjacent and flipped spatially. In that case, the pickup-induced voltages should cancel while the current-induced voltages should add, as desired. It's hard to find CMRR values for high-end humbuckers, especially since they are usually intentially imbalanced for tone considerations, but I would venture a guess that the ~40-50 dB required in this case is not completely out of the question. I've ordered another identical inductor to see what we can do.

  1236   Wed Apr 8 21:12:19 2015 DmassCryostatSiFiRe-evacuated large cryostat

You can make them yourselves if you want with activated charcoal and stycast - you just need to bake and pump to clean it. I got some activated charcoal from Keith a while back (it's either in the glass cabinets in the Cryolab, or the fume hood in the ATF), but I think you'll need to buy some new stycast (which you should do anyways if you're going to be gluing stuff in there - the stuff we had for the Cryolab is past its expiration). The cans of what you want are in the ATF fume hood.

If you have room somewhere by a screw hole, you just make machine a mounting block, cover it in HV compatible epoxy (stycast), and roll it around in activated charcoal. To clean it, just pump and bake (the setup to do this for this should be in the CTN lab, if you cannot figure out what it is / how to use it and are nice to him, Dmass will probably help you). If you do it this way, you can unscrew it and pump/bake it when it loses its vacuumy goodness/cleanliness.

If you don't want to keep vacuum but don't want to ruin your getter, just backfill with N2 gas the valve off. Keeping the cryostat open will unavoidably dirty it up though.

  1235   Tue Apr 7 16:24:18 2015 ZachCryostatSiFiGetter quote

Apparently, the getter is made by IRLabs themselves. They've quoted me $690 for a replacement (see attachment *removed since I was informed that's not allowed*).

Quote:

Good point. If you or Nic can find it, please post some getter info and we can get a spare just in case this one is already stuffed.

 

  1234   Sat Apr 4 18:08:18 2015 ranaCryostatSiFiRe-evacuated large cryostat

Good point. If you or Nic can find it, please post some getter info and we can get a spare just in case this one is already stuffed.

Quote:

I roughed our cryostat back down tonight. Remember, this model has a getter that degrades over time when exposed to air, so we should minimize the time that it's not under vacuum.

 

  1233   Thu Apr 2 20:21:18 2015 ZachCryostatSiFiRe-evacuated large cryostat

I roughed our cryostat back down tonight. Remember, this model has a getter that degrades over time when exposed to air, so we should minimize the time that it's not under vacuum.

  1232   Thu Apr 2 01:03:33 2015 ZachLab InfrastructureSiFiRack set up, table cleared and real layout started

Today, I installed the shelves into our rack and moved the electronics that were on the table into it. I then cleared the table and started setting things up the way they will be.

The cryostat is now on the south side of the table, and the lasers are set up on the north side. I connected and placed the AM and PM modulators in both laser chains, then placed the output couplers. I used posts to protect the fibers and strain releif the SMA cables going to the modulators, and installed the BNC patch bay on the end of the table near the rack. I'm waiting on our 3/4" posts to continue with the optical layout.

With everything connected, I trimmed the AM DC voltages to maximize the transmission and verified that the output power for each was ~25% what it is directly out of the laser, as it should be (since each modulator has loss ~3 dB).

 

  1231   Mon Mar 30 17:56:13 2015 DmassNoise HuntingNoise BudgetWest Photothermal Transfer Function Actually Measured

Loss vs T. Confused on error bars here, so ommitted for now. This will have error bars before I thesify it.

I used the transfer function as well as our knowledge of the noises to estimate the coherence of each meeasurement, and then that to estimate the fractional uncertainty of the (SR785's estimate of the) transfer function, but the fractional error was significantly smaller than the residuals (e.g. the systematic contributions to the reduced Chi^2 dominate over the random noise).

Error bars can still be derived from the curvature of the residuals over the fit parameters.

I tried to use MATLABs curve fitting package to give me error bars from variance over the measurement, but couldn't figure out how to give it the farsi code results in a way that it would fit over.

 

I spent a while trying to measure the PT TF @ 125K and 114K, but:

  • The 125K measurement point is all noise for unknown reasons (see elog:1226 for raw data)
  • The 114K measurement baffles me - I think the optical offset coupling (flat) and the PLL offset sensitivity are crossing, but they cross with a slope difference of f (f^0 and f^1), and notch. I thought about it for a while but do not know what could cause this phase disparity.

If I spend a while (~half day? day?_ longer on fitting with the 114K measurement, I might be able to add a point to the plot. Time is precious right now so I am tabling it.

Given the uncertainty in why I couldn't measure the TF cold, I am tabling the idea of having another cooldown to get more data until after I defend. The best scientific statement I have right now is:

"The results indicate that further exploration of the transfer function is warranted*"

*if anyone cares about the results.

 

I took cavity pole measurements (via transfer function)

Attachment 1: lossvsT.pdf
lossvsT.pdf
  1230   Fri Mar 27 00:26:02 2015 ZachSummarySiFi - ringdownTaiwan cantilever phi vs. T (old but unreported)

Before the LVC meeting, I had just done a long steady-state Q measurement on the Taiwan cantilever. I got too distracted by the melted epoxy disaster (CRYO:1225) to actually post the data.

Below is a plot of the ~16 hour stretch of data (second trend), showing the temperature and instantaneous loss angle. The temperature was stepped in 10-K increments from 90 to 130 K, holding at each temperature for 3 hrs to allow the system to equilibrate and integrate (except for some of the early steps which required some manual intervention).

The main result is that the loss seems relatively constant at ~10-6 from low temperature to ~120 K, where it starts to increase. Towards the end of the 130-K stretch, the LN2 ran out, and the system started heating to room temperature uncontrolled (i.e., heater output was railed at zero).

This level is too high to be from the Si, so I assumed it was some residual clamping loss. I was dubious that the figure from the one reference that Matt A. found and gave me for the cryogenic Q of stainless steel would be applicable to our particular clamp, so I thought I might try to measure it directly, in parallel with the cleanup of mess in the cryostat. To do this, I got some spare steel wire from Gabriele and made a makeshift suspension, hanging the top piece of the clamp, hoping to measure the loss of its lowest vibrational mode. I knew it was a long shot, since this mode should be around 17 kHz, but I set it up in the simple vacuum chamber anyway, and tried to excite it and read it out optically. The first bending mode should have nodes *near* the suspension points, so I thought I might get some kind of meaningful results if I could actually see a ringdown.

I was unsuccessful. I tried various excitation schemes, from broadband (banging stuff) to narrowband (bandpassed white noise, amplified with the boom box and blasted out of a speaker touching the chamber), and none revealed any mode excitation. I was able to see broadband noise increase with the excitation profile, but no lines, so most likely I was seeing some alternate path.

I still think it would be nice to get an empirical measurement of the cryogenic Q of the steel we use for our clamps. Maybe we can set up a laser vibrometer measurement like Norna and her student did a few years back on the steel gyro PMC?

  1229   Thu Mar 26 20:06:02 2015 ZachCryostatSiFi - ringdownSmall cryostat reassembled, Taiwan cantilever in clamp, pumping down

[Den, Chris, Nic, Zach]

Since my snafu before the LVC meeting (CRYO:1225), the small cryostat has been in pieces being thoroughly cleaned and aired out. Nic wanted to have the ringdown setup rebuilt so that we can demo the steady-state Q measurement technique for our visitors, so we did some work today to make that happen.

This morning, I re-lined the main chamber walls and floor with aluminum tape. This model came with some thin foil lining the walls, attached by periodic thin strips of double-sided paper tape. We have been intermittently scraping some foil off each time we cycle, and since a nasty residue was present on the floor of the chamber after the epoxy incident, I figured it was time to replace the lining. I just used aluminum tape since a.) it is stronger and will be less prone to scraping off, and b.) if and when we need to replace it again, it should come off much more easily.

This afternoon, we rebuilt the cryo package on the cold plate (clamp with Taiwan cantilever installed, ESD, and 45º mirror). Since we don't want to use epoxy to mount the power resistor anymore and we don't have any tapped holes in the clamp, we have not equipped any heat source or temperature sensor. This is fine, since we really just want to use it as a demo this time around, and room temperature should be sufficient. If we want, we can still cool it down to LN2 temperature, but we won't have any actuation or readout.

Upon pumpdown, we noticed that the pressure had stalled at around 20 mTorr after a good 20 mins of pumping, indicating that we had a leak. We checked the top seal and electrical feedthrough (which had also been freshly reattached during the rebuild), and found no issues. With nothing else to try, we decided it was most likely the seal between the chamber floor and the main section (I had to foil this with rectangular sections of tape, which I then XActo cut into a circle at the o-ring groove, so it was possible that a foil flake was blocking the seal). With everything still in place, we flipped the cryostat over and removed the bottom. We found a couple places where a tiny piece may have extended into the seal, so I re-cut the circle more conservatively. When re re-sealed, we found the pumpdown profile to be much closer to what we usually expect. The pressure was a few mTorr after ~10 minutes and showed signs of healthy decline.

We rebuilt the optical readout, then tested the MODERINGER amplitude sensing and found everything seemed to be working. We did not want to test the ESD at this high pressure. When I left, the Q was relatively low at maybe a few thousand, but gas damping was likely still a limiting factor. Also likely is that there is still some residue on the cantilever that I didn't get off, or perhaps even that some irreparable damage might have been done. We should be able to tell when the pressure is low enough.

  1228   Tue Mar 24 14:34:26 2015 ranaThings to BuyVacuumCryostat unpacked (x-post from SUS elog)

I opened one of the viewports today to inspect it for window size. It does indeed look like the AccuGlass 2.75 CF (there are types of 2.75 CF on the web, so no real standard).

The window itself is a nonstandard size. It is pressed down on top by a aluminum ring and the seal is made on the bottom by a Viton O-ring.

Windows are on order - expected to be here by next week. T = 99% @ 1550 nm. I don't think there's an intentional wedge.

 

Attachment 1: AccuGlass275CF.pdf
AccuGlass275CF.pdf
Attachment 2: ag_arcoating_graph_web.jpg
ag_arcoating_graph_web.jpg
  1227   Mon Mar 23 14:23:09 2015 DmassNoise HuntingNoise BudgetWest Photothermal Transfer Function Actually Measured

Ran the modified Farsi code with temperature dependance added in for:

  • Silicon Thermal Conductivity
  • Silicon Thermal Expansion
  • Silicon Specific Heat

 

Attachment 1: SubstrateThermalvsT.png
SubstrateThermalvsT.png
Attachment 2: ThermoOpticvsT.png
ThermoOpticvsT.png
Attachment 3: CoatingThermalvsT.png
CoatingThermalvsT.png
Attachment 4: CoherentSumvsT.png
CoherentSumvsT.png
Attachment 5: PthermBudget114K.png
PthermBudget114K.png
Attachment 6: PthermBudget125K.png
PthermBudget125K.png
Attachment 7: PthermBudget200K.png
PthermBudget200K.png
Attachment 8: PthermBudget300K.png
PthermBudget300K.png
  1226   Thu Mar 19 14:00:21 2015 DmassNoise HuntingNoise BudgetWest Photothermal Transfer Function Actually Measured

Raw Data from photothermal TFs:


Calibrations in parentheticals.

114K calibration is different, but not far off (A) and (B).

114K response consistent with PDH optical offset + PLL offset induced RIN -> phase sensitivity. Could fit these two effects and see if what was left looked like a real measurement of photothermal response

Am not sure why 125K measurement was so noisy, but hacked at it for a while and couldn't get anything that looked all that coherent out of it.

Warmed up slowly, turning on and off heaters to get @ data points:

No points between 125K and 233K because Cryostat temp ran away very rapidly overnight after a couple days of very slow warming and patient watching :(

All photothermal fits are generated using material properties at 300K, so caution should be taken interpreting these results (e.g. using the fit of the room temperature model to the cryo data introduces unknown frequency and temperature dependent systematic errors into the fit / estimation of loss).

Can use beat noise, and bandwidth to generate SNR(f), and averaging time to turn that into frequency dependent error bars if we wanted to generate good error bars on the 300K loss number. Unsure how valid it would be to do this on other points because of the known unknowns (systematics) in the model.

Can

Attachment 1: PthermDataRaw.pdf
PthermDataRaw.pdf
Attachment 2: PthermDataRaw.pdf
PthermDataRaw.pdf
Attachment 3: Pthermfit300K.pdf
Pthermfit300K.pdf
Attachment 4: Pthermfit290K.pdf
Pthermfit290K.pdf
Attachment 5: Pthermfit283K.pdf
Pthermfit283K.pdf
Attachment 6: Pthermfit255K.pdf
Pthermfit255K.pdf
Attachment 7: Pthermfit233K.pdf
Pthermfit233K.pdf
Attachment 8: Pthermfit233K.pdf
Pthermfit233K.pdf
  1225   Thu Mar 12 18:45:50 2015 ZachCryostatstuff happensEpoxy mess

I brought the cyrostat up to room temperature today to do some work inside, and there appears to have been a minor disaster.

The 100-Ohm heater was was glued to the large steel clamping block with (I'm assuming) conductive epoxy, and it appears as though there was enough heat to dissociate it at some point during the latest run.

When I opened the cryostat, there was a strong noxious smell, which was my first indication that anything had gone wrong. Upon opening the top of the can, I found that the power resistor was now attached to it and not the block (the whole thing hangs upside down, so the resistor fell of, then grabbed onto the lid as it cooled).

It's unclear when this happened. I had been running the cantilever at constant amplitude over varying temperature over the last 2-3 days and there was no significan event to indicate the time when the resistor must have fallen. My best guess was that it happened as I provided a steady power of ~10 W to it this morning to speed up the heat-up. The RTD, located on an adjacent side of the block some ~3 cm away, never registered above room temperature, and that was at the very end before I vented the chamber.

Thankfully, this stuff dissolves readily in isopropanol. I've been going through the entire system, starting with the Taiwan cantilever itself, and cleaning everything. There is a very slight (and typically invisible) residue apparent on most parts of the chamber, so I'm getting that all off.

The power resistor on the smaller (radial) clamp is screwed on, which I think is a better solution in light of this.

  1224   Tue Mar 10 03:51:42 2015 ZachDailyProgressSiFi - ringdownTaiwan cantilever long-term low-temperature phi

I reported in the replied-to entry that the Q of the Taiwanese cantilever at low temperatures actually appeared to have gotten lower at low temperatures, relative to the case before the Si spacer was added to the clamp to avoid skewness. However, the data from the longer run over this past weekend (see the ~20-hr stretch below) seem to suggest a Q not significantly different from that measured in CRYO:1213.

Interestingly, the online phi measurement starts out at the higher level I indicated in the previous post, but then slowly approaches a level not inconsistent with the ~1.5 x 10-6 number from before the spacer addition. The title is misleading, as the system actually approached a minimum temperature of ~90 K on Saturday, but the thermoelastic noise prediction is roughly flat over this temperature band, so that shouldn't be a factor, and the associated deflection from this temperature shift should not be enough to account for this drift via calibration error.

As I discuss in the quote, I had hoped to make a continuous phi measurement as the system warmed leading up to today, but at the time I neglected to consider the thermal deflection, which over such a large temperature swing completely rasters the beam off the QPD. In retrospect, I'm lucky that this effect didn't break the cantilever as the sensing gain was reduced from the misalignment---thankfully, the loop destabilized quickly enough that the watchdog script killed the feedback before anything happened.

So, it looks like we'll have to make this measurement the old fashioned way, point-by-point, which is why I spent time reconfiguring the temperature control today. I'm running an active measurement overnight at 120 K to see if we see a Q bump there.

Quote:

As you can see, the loss is hovering around 3 x 10-6, giving a Q around 3 x 105, which is slightly but significantly lower than what we measured before I added the Si spacer to avoid skewing the clamp (CRYO:1213). I would chalk this up to the spacer actually making the clamp worse, but we did in fact see a huge improvement at room temperature (CRYO:1216). So, like, what the hell man?

I've left it running to collect more data over the weekend. I haven't gone over the temperature readout/control system with Nic, so I set up a simple temperature readout in the meantime so that we can have at least a coarse Q(T) measurement as it warms. To do this, I simply put a 1k resistor inline with the RTD and put 5V across with a lab supply. The second set of RTD leads goes to the temperature readout input in the digital system, so this is now just a DC readout of the voltage across the RTD. The lockin input channel X1:SCQ-TEMPERATURE_LOCKIN_DEMOD_SIG_OUT is calibrated to volts, and is equal to 5 V * RRTD/(1k + RRTD).

 

  1223   Tue Mar 10 03:19:49 2015 ZachComputingSiFi - ringdownTemperature calibration and control changes

It was a little unclear to me how the digital temperature control was supposed to work as it was built, so I made some modifications today.

Readout / Calibration

The previous implementation used a digital lockin setup (as does the new one), but the output of this was converted to an error signal for the temperature control loop using a relatively primitive calibration, so I added some math into the model to make it a little more exact. The changes can be divided into two sections: 1) the demod voltage to RTD resistance section and 2) the RTD resistance to temperature section.

Demod voltage to RTD resistance

To get a nice linear temperature signal using the 4-lead sensing method, the RTD current should not be determined solely by the RTD (otherwise, the readout just sees the excitation directly). So, I have added a 1-kOhm resistor in series with the RTD in the LO path, just as I did with the temporary setup described in CRYO:1217. The series resistor and the RTD now form a voltage divider where the RTD resistance can be inferred to high precision in the limit RRTD << Ri (= 1 kOhm). This is almost always true, but the model does include the exact expression for RRTD(Vout). To set the calibration, one must enter 2 values:

  • RISET (aka Ri): The input series resistor (1 kOhm presently)
  • VLO: The (pk) amplitude of the LO signal in volts across Ri+RRTD

RTD resistance to temperature

For this, I have implemented the full, quartic formula from the ASTM standard (see our RTD manufacturer's page). Other than 4 preprogrammed empirical constants (and the Kelvin conversion of +273.15 at the end), this only requires one input from the user:

  • R0: The RTD resistance at 0° C (100 Ohms for our RTDs).

 

Heater actuation

The heater actuation section was in pretty good shape, so I didn't really have to make any modifications there. One thing I did do was add a heater power calculation, which requires the user to enter the heater resistance.

On the hardware end, since I'm using the bigger steel block clamp which also has the higher-resistance (100-Ohm) power resistor, I found that I needed more juice than what the DAC -> voltage amp/buffer that Marie and Nic used could provide (this circuit was regulated to 15 V, giving a max power of 2.25 W, which just isn't enough). I stole my Sorensen HV supply back from the CTN lab for now, as it seems to be unused, and used it as a HV amplifier via the external control feature. Since this unit doesn't allow voltage range limiting in remote mode, I added a 1/10 divider between the DAC and it so that I didn't have to trust software limiters. Really, I should attenuate the HV output, but I couldn't think of an easy way to do that with the stuff I had on hand. Anyway, the railed heater voltage from a +10 V DAC signal is ~40 V --> 16 W.

 

Finally, I edited the SCQ master screen to reflect all these changes. Here, you can see the system being held at 120 K:

 

  1222   Sun Mar 8 18:14:39 2015 DmassNotesCavitySubtraction

The main thing I wanted to do with the coherence+ beat @ 300K beat data is subtraction.

  • Equation 7.33 of Bendat & Piersol applies here since Coh(PDH_E,PDH_W) = 0:
    • PSD(resid.noise) = (1-Coh(PDH_E,PLL)-Coh(PDH_W,PLL))*PSD(PLL)
    • aka ASD(resid.noise) = sqrt(1-Coh(E,PLL)-Coh(E,PLL))*ASD(PLL)
    • aka subtraction level = sqrt(1-Coh(E,PLL)-Coh(E,PLL))

Something sort of interesting (but not unexpected when you think about it) happens when you start to combine multiple coherences in this way:

If the (quadrature) sum of errors in your estimates of coherence (mag^2(rmsav(CSD))/(PDS1*PSD2) in this case) is sizable compared to the residual noise, equation 7.33 can give you negative values of power (which corresponds to imaginary values of ASD).

I pondered for a bit and came to it is very unsurprising that subtracting two things of similar value (1, sum of coherences) to get to a smaller thing (tiny residual noise) comes with its own problems.

Here is the subtraction residual:

Multiply the calibrated beat signal by this to get the answer "what is our best upper bound on the residual noise in the beat not due to unsuppressed frequency noise?"

The 1kHz to 10kHz slope is where the PDH loops shift from being noise limited to gain limited.

  1221   Sun Mar 8 17:18:52 2015 DmassNotesCavityCoherences

Coherence measurements:

 

The SR785 estimates coherence as follows:

  • Coh(A,B) = Mag^2(rmsavg(CSD(A,B)) / (PSD(A)*PSD(B))
  • From the wording in the SR785 User's manual, the PSD averaging is set by the measurement type, and each measurement of the CSD is RMS averaged. I believe this means: as long as we are RMS averaging, everything will be in RMS.

We can use the coherence measurements to tell us what PDH_E/PDH_W is, which can help illuminate the earlier "what's up with the PDH calibration" problem.

sqrt(Coh(PDH_E,PLL)/Coh(PDH_W,PLL)) = ASD(PDH_E)/ASD(PDH_W)

What's going on?

  • Maybe we wrote down the file names wrong and switched them?
    • The yellow trace shoots this theory down - shape of West/East cannot produce the shape of coherences we measured
  • Maybe the PDH loop was set differently between the coherence measurements, the beat measurement, and the PDH error signal measurements
    • The flatness of the bottom traces indicates that the spectral shape of E/W was consistent between the Coherence measurements and the PDH error signal measurements
  • Maybe the DC numbers for the PDH calibration I used are just wrong
    • This seems to be the case

PDH_E[Hz] / PDH_W[Hz] * 0.88 = PDH_E[V] / PDH_W[V], so our PDH calibrations have to reflect this ratio:

EastCal [Hz/V] / WestCal [Hz/V] = 0.88

The ratio of the calibrations I used in elog:1219 was incorrect by a factor of 2.1.

Restricting the ratio of Ecal/Wcal to be 0.88, and tuning the W PDH cal up to 5e5 Hz/V, I get:

The DC calibrations quoted in elot:1219 are incorrect for unknown reasons (My money is on "oops I forgot that I added an attenuator" error which makes the demod gain I used (0.63) incorrect.)

  1220   Sun Mar 8 16:04:23 2015 DmassNotesCavity300K Transmitted Spectra from Dec 2014 run

Transmissed RIN spectra

Trans noise level was minimized by tuning the optical PDH offset (generated via RFAM, tuned via waveplates by the EOM), and the electronic offset (tuned via low noise offset adjust knob on LB1005 box)

Measured with PDA20CS on pickoffs from each path. Trans DC voltage ~1V, TransZ gain 10 or 20 dB (not PDA20CS dark noise limited)

  1219   Sun Mar 8 15:20:10 2015 DmassNotesCavity300K PDH data from Dec run

PDH error signal measurement:

Measured using the "error monitor" output of the LB1005 boxes

High frequency portion is multiplied by 2 b/c 50 ohm impedance of 4395 + LB err mon

 

PDH OLTF's:

OLTFs taken by measuring between two points on the loop, injecting on either side of the points, then dividing out those measurements - the extra measurement saves on calibration time sunk.

Calibration:

  • [V/Hz]_E = (2*Pinput*gamma*sqrt(coupling)/fcav) * transZ * demodgain = 5.72e-6 V/Hz (aka 1.75e5 Hz/V)
  • [V/Hz]_W = 2.7e-6 V/Hz (aka 3.7e5 Hz/V)
  • The difference between the two calibrations is consistent with what we know about the tow paths: East had a lower cavity pole, higher transimpedance gain, and higher gamma (e.g it's not ann agebraic error)

Calibrated PDH err plot vs PLL:

Something is clearly wrong with the calibrations I used for the PDH error signals. I suspect the PDH East calibration based on morphology (when I increase the E PDH calibration by ~5x, the quadrature sum of E and W error signal residuals line up with the calibrated beat signal, which is consistent with the sum of coherences being near unity (coherence measurements between PLL, PDH_W, and PDH_E indicate that unsupressed laser frequency noise is 10x higher than any other noise source at frequencies between 10kHz and 100kHz [no data on coherence above 100kHz])


The PDH calibrations can be verified/overthrown by using the measured PDH OLTFs and dividing out the known parts of the plant.

  1218   Sun Mar 8 14:37:18 2015 DmassNotesCavity300K beat data from Dec run

Beat raw ASD from 300K Dec measurement set:

The PLL setup was a Marconi and an SR560 after the demod/lowpass.

  • State A: Marconi range = 10 kHz  // 560 gain = 20
  • State B: Marconi range = 40 khz // 560 gain = 10
  • Range was set so high because cavity drift is high at room temperature (\alpha (T) ~10^-6 @ 300K for Si)

PLL noise was determined by replacing the transmitted beat with another marconi at the same RF power / frequency, with the VCO functionality disabled (the noise goes WAY down when you do this). Input noise of marconi we were using as a VCO was the dominant effect at high range level. Range was set as low as seemed reasonable to trade off between averaging and noise. The bucket noise is definitely not noise from the PLL readout.

PLL Calibration:

OLTF was measured as G/(1-G) in each state

Control signal calibration is V_ctrl*(marconirange/1.41V)*(1-G)/G

  • Marconi range is given in mixed units (pk/rms) so we divide by 1.41 to give it in rms/rms

Error signal calibration is  V_err/(7e-2 V/rad)*(1-G)/1.41

  • 7e-2 V/rad determined assuming that the V/rad disciminant response is white in amplitude past the UGF of the PLL, then dividing out the 560 gain and marconi response from the OLTF measurement. Both state A and state B gave the same number
  • extra 1.41 at the and of error calibration is a 50 ohm thing

Calibrated error and control signals for the beat are:

Data/plots/matlab code for processing all on svn in CryoLab/Measurements/CoherentSub/ directory

  1217   Fri Mar 6 19:02:28 2015 ZachDailyProgressSiFi - ringdownQ at low temperature after reclamp… worse?

The cantilever had cooled to around 100 K by this morning, so I set up the mode ringer and began an active measurement on the fundamental mode. The online loss angle measurement for a 3-hr period beginning around an hour after lock is shown below (this is the control signal filtered by a 2nd-order low pass at 0.2 mHz.

As you can see, the loss is hovering around 3 x 10-6, giving a Q around 3 x 105, which is slightly but significantly lower than what we measured before I added the Si spacer to avoid skewing the clamp (CRYO:1213). I would chalk this up to the spacer actually making the clamp worse, but we did in fact see a huge improvement at room temperature (CRYO:1216). So, like, what the hell man?

I've left it running to collect more data over the weekend. I haven't gone over the temperature readout/control system with Nic, so I set up a simple temperature readout in the meantime so that we can have at least a coarse Q(T) measurement as it warms. To do this, I simply put a 1k resistor inline with the RTD and put 5V across with a lab supply. The second set of RTD leads goes to the temperature readout input in the digital system, so this is now just a DC readout of the voltage across the RTD. The lockin input channel X1:SCQ-TEMPERATURE_LOCKIN_DEMOD_SIG_OUT is calibrated to volts, and is equal to 5 V * RRTD/(1k + RRTD).

Quote:

The LN2 dewar was refilled today, so I filled the cryostat and we'll see how it looks at low temperature tomorrow.

 

 

 

  1216   Thu Mar 5 22:35:43 2015 ZachDailyProgressSiFi - ringdownTaiwan room-temp Q > 10^5, cooling now

On Tuesday night, when I added the Si spacer to the clamp, I measured a Q of ~46000, but I noted that it had been increasing up to that point, likely due to the residual gas damping (see CRYO:1214). Last night, I made another measurement and found it to be much higher, at ~1.2 x 105 (tau ~ 350 s). This is much better than we have seen at room temperature thus far, so it looks like my spacer addition has helped.

I remeasured this an hour or so later and saw no appreciable increase. I checked again today and it appears as though it may have increased slightly, but it was hard to say for sure due to higher environmental noise. Really, we need the steady-state ringdown to make a good measurement at this level.

The LN2 dewar was refilled today, so I filled the cryostat and we'll see how it looks at low temperature tomorrow.

  1215   Thu Mar 5 22:20:06 2015 ZachLab InfrastructureCryoLN2 refilled, new dewar #102

I called the campus service yesterday morning to have the LN2 dewar refilled. They got around to it today. New dewar number is 102.

  1214   Wed Mar 4 02:32:45 2015 ZachDailyProgressSiFi - ringdownSi spacer added to clamp holding Taiwan cantilever

The most recent measurements on the Taiwan-sourced Glasgow-style cantilver (see CRYO:1213) are encouraging, but the best Q measurement at low temperature is still a couple orders of magnitude worse than what is theoretically achievable, and about one order of magnitude worse than our conservative clamp loss estimates. Also, I've done some measurements on other modes (that have different expected clamp loss contributions due to the relative strain energy ratios) to try and sort out what is going on, with little success. Finally, some modes---including the 2nd bending mode at ~650 Hz---exhibited very low Q for no known reason.

One thing I thought about is that, since the Taiwan cantilever did not fit in the groove that was built into the block for the Glasgow-style cantilevers and therefore is just sandwiched between the two large pieces making up the clamp (see CRYO:1211), the clamp is likely pushing down at somewhat of an angle, which could lead to all sorts of non-idealities. Since the other Si samples we have lying around are roughly the size of the clamping region of this cantilever (~300-500 um), I opened up the cryostat today and reclamped the cantilever using a spare broken-off 300-um-thick cantilever piece as a spacer on the other side:

Pumping it all back down, I immediately measured Qs a bit higher than what we saw last time around at room temperature. The last measurement I made before leaving was tau ~ 135 s ==> Q ~ 46000, though it had been increasing up to that point, likely from the residual pressure, which was at ~10-3 Torr when I left. Compare this with the Q of ~14000 from the last time around, though admittedly I did not record the pressure at which this was measured.

  1213   Fri Feb 27 05:47:27 2015 ZachDailyProgressSiFi - ringdownTaiwan cantilever fundamental mode Q

[Nic, Zach]

We measured the Q of the fundamental (~106 Hz) mode of the Taiwan cantilever in two ways. First, we used Nic's active steady-state method, and then we did a traditional ringdown. The results seem to agree, but the precision of the first method is much better due to the dynamic range of the readout for this mode: the motion becomes nonlinear at an amplitude only a few times greater than the background excitation level. Over a ~4-hr average, the loss is measured to be 1.45 x 10-6 ± 2.9 x 10-7, giving a Q of ~6.9 x 105.

Here is a plot of the instantaneous phi from the calibrated control signal. This data has already been fed through a ~1-hr lowpass, and then the data from the initial settling time has been truncated away. The mean and standard deviation of the rest of the points are what is reported.

After this measurement was made, we shut off the servo and allowed the mode to ring down. Here is that ringdown, along with a predicted range of theoretical curves using the result from above. As you can see, they are fairly consistent with what is measured, considering that the system quickly reaches a regime where it is excited by the environment (that is, only the initial part of the ringdown, where the agreement is good, is very trustworthy).

This Q is a couple orders of magnitude lower than what is expected for this mode at this temperature, but it is also only a factor of 2-3 worse than the best measurements using a similar apparatus at Glasgow (to my knowledge).

It bugs me that we don't seem to have any information about what steel looks like at low temperatures. Given my COMSOL strain energy modeling, the energy ratio for this mode is about 3 x 10-4, so this could be explained by clamp loss if the steel Q is as low as a few hundred. I'm looking into other modes to try and support or refute this hypothesis; since different modes have different energy ratios, we may be able to see what's going on. In parallel, I'm asking Matt and others to find out what is really known about cryogenic steel.

 

  1212   Thu Feb 26 18:47:59 2015 ZachLaserM2 ISS3-mm diode dark current and noise

Given that we see some excess noise in our 3-mm Laser Components diodes (IG17X3000G1i), especially with one of them, I went ahead and did a more careful measurement of both the dark currents and noise.

To make this measurement, I switched to OPA140 transimpedance amps on the M2 readout board and used a 1-M\Omega transimpedance.

Dark current

The OPA140 has a bias/offset current of 10 pA max and an offset voltage of 120 uV max, the latter which therefore limits the DC current sensitivity with this transimpedance to 0.12 nA. There is also some allowed current variation over temperature (±3 nA bias and ±1 nA offset over -40 to +125 °C), so this can add some more DC uncertainty if the lab temperature is a few degrees away from 25 °C. Plugging the outputs into the DMM without the diodes connected, I measured 0.0 mV and 0.2 mV on amps 1 and 2, respectively. This is consistent with the op amp spec.

The Laser Components spec for the dark current (at 5-V bias, where I measured it) is 20 nA typ, 100 nA max. Plugging in the diodes while keeping them in a blocked box and with the room lights off, I measured the following bias currents (output voltage divided by 1 M\Omega):

  • S/N 7842: 78.9 nA
  • S/N 7843: 61.7 nA
  • S/N 7844: 43.8 nA
  • S/N 7845: > 200 nA (started below 100 nA, increased continuously while energized---see below)

The first 3 diodes are within the max spec, while 7845 seems to be exhibiting some catastrophic failure mode where the dark current is avalanching whenever the bias is engaged. Below is a plot of the measured amplifier output after a turn-on of this diode, with one of a healthy diode for comparison to the right. This was taken in the middle of the testing, and the last measurement of the current before this turn-on was around 140 nA. As you can see, there is an initial slew (not inconsistent with the timescale of the bias turn-on), followed by a slow but monotonic increase of the dark current over time. When this was repeated, the initial slew brought the current again to the last-known highest level.

 

So, as you can see, S/N 7845 is clearly broken. Maybe we can get a replacement

Noise

I used the same transimpedance amp setup to measure the noise. All diodes show spectactularly higher noise than the advertised level of 3.2 x 10-14 W/rtHz NEP (~3 x 10-14 A/rtHz), with a 1/f characteristic that, if extrapolated, would not intercept the quoted spec until ~1 MHz. A frequency for this spec is not mentioned on the datasheet. In all cases, the circuit was allowed to equilibrate for a few minutes before a measurement was made. The spectra below were found to be stationary, with the exception of occasional glitches.

The readout noise is limited from several 100 mHz up to near 1 kHz by the Johnson noise of the 1-M\Omega transimpedance resistor, above which there is some noise peaking centered around 40 kHz that is not inconsistent with other measurements I have made with this op amp in very-high-impedance environments (c.f., 40m:8151).

What gives?

Quote:

IF the DC dark current is out of spec, we might be able to get a replacement. Might be specs on the website. I think Frank had a Keithley instrument to measure dark currents that are low - probably in his diode destruction elogs or DCC docs.

 

  1211   Wed Feb 25 04:29:49 2015 ZachDailyProgressSiFi - ringdownTaiwan cantilever has higher Q, going for cryo cycle now

[Nic, Zach]

Nic got a Glasgow-style cantilever from a group in Taiwan, and a quick test in the rapid cycle chamber showed that it had pretty low loss, so we are running it in the cryostat now. As a reminder, these are the rough dimensions of this style cantilever:

Below is a photo of the box it came in, showing the actual 92-um thickness of this sample, as well as a shot of it in the vacuum chamber. For some reason, this particular sample's clamping tab did not fit in the groove that Nic had built into the clamping block for the other Glasgow cantilevers, so I had to mount it to the side against the flat faces of the clamp (as I've been doing with our larger samples).

 

This evening, we transferred it over to the cryostat and restored all the electrical connections for what will hopefully be a fruitful cryo run. Here is a ringdown of the fundamental mode (~106 Hz) at room temperature:

The measured decay time of 41 seconds corresponds to a Q of around 14,000, which is about as good as we expect at room temperature. This sample is probably better than our other ones for at least 2 reasons:

  1. It is made from a better-quality (FZ) wafer, and
  2. It has been manufactured monolithically with a thicker clamping tab, which our modeling suggests is a very effective way to evade clamping losses by keeping strain energy within the silicon.

Given that we didn't see much improvement at all with our other samples when going to low temperature, I believe (2) is by far the biggest effect. The Glasgow wafers only have the clamp-region thickness extended to one side, which is modelled to be worse than if you go both ways, but it is still much better than we can do with our discrete sandwiching.

I filled the LN2 reservoir and the volume is cooling overnight. I did some rough ringdowns at a point when the steel block was registering around 160 K and found greatly improved Qs already (approaching and perhaps exceeding 105). We will continue to make measurements tomorrow.

  1210   Tue Feb 24 05:17:04 2015 ZachDailyProgressM2 ISSFirst real M2 test

Tonight I succeeded in using the M2 ISS readout board and the 3-mm diodes to do some real intensity stabilization using the SiFi test setup.

First, I built a foam box to use as a temporary enclosure for the PMC and diodes until we get our real box finished:

There are holes for the input and REFL beams, and the diodes are held with makeshift mounts that clamp down on the sockets. Clearly, these aren't as stiff or stable as what we're having built, but they do the job for now. There is a steering mirror before the 50/50 BS so that the position on each diode can be adjusted separately with ease.

I didn't want to open the Chachi ISS box can of worms yet, so I just built my own temporary breadboard circuit. I had done some preliminary SR560 locking, so I knew roughly what I wanted, and I measured the modulator -> PD transfer function again today and verified that it was flat well above 100 kHz. I made a 2-stage pole/zero-style circuit, with a double (removable) pole at 300 Hz and a zero at 10 kHz to bring phase back to -90° around the target UGF of 100 kHz. It looks like this:

I wanted to DC couple, so I came up with an idea to pick off the stable 5-V bias supply from the M2 board and sum it with the (negative) output of the in-loop PD in the first stage of the servo. I had some current-related issues with the summer at first, but these went away when I increased the input resistors a bit (n.b., to fix the gain I had to change some other components, and as a result the controller TF is actually slightly different than shown above, but not much).

Hooking it up, it locked right away with the expected UGF of near 100 kHz (not yet measured, but inferred from the transfer functions and spectra). Here is the stabilization result:

As you can see, the out-of-loop signal is stabilized to the shot noise level (which is \sqrt{2} higher than the bare shot noise for half the beam due to the well-understood correlated noise imprinted by the loop) from about 5 kHz down to just below 100 Hz. Below this, there is clearly some differential environmental noise between the PDs. I did some beam scanning to try and minimize with some success, but not much. I'm not sure what the coherence below 20 Hz indicates---the in-loop signal is suppressed to below the measurement noise level, while the OOL signal exhibits excess differential noise, so I don't see why there should be any coherence.

In any case, this is a nice verification that:

  • The M2 readout board works with real optical signals
  • The intensity feedback system for the SiFi experiment works
  • The 3-mm diodes (save for the one bad one---see CRYO:1207) behave nicely, at least for these relatively low powers
  • The PMC -> PD scheme shows promise for our future tests with nicer hardware
  1209   Mon Feb 23 19:55:08 2015 KojiLaserM2 ISS3-mm diode initial noise measurement

Frank had = PeterK had = It went to LHO

I wonder if it helps to use FEMTO's DLPCA-200 that I have somewhere in my lab.

  1208   Mon Feb 23 17:57:49 2015 ranaLaserM2 ISS3-mm diode initial noise measurement

IF the DC dark current is out of spec, we might be able to get a replacement. Might be specs on the website. I think Frank had a Keithley instrument to measure dark currents that are low - probably in his diode destruction elogs or DCC docs.

  1207   Mon Feb 23 14:57:23 2015 ZachLaserM2 ISS3-mm diode initial noise measurement

I was preparing to do an initial test of the M2 ISS readout board with the 3-mm diodes on the SiFi test setup when I noticed some anomalously high noise on one of the diodes. So, I decided to make a more careful measurement and test all 4 diodes. I found that only one (S/N 7845) exhibits this very bad excess 1/f noise, but all four have it present at some level.

For this test, I had the transimpedance fairly high at Z = 2.7 k\Omega since I am only working with < 5 mW of power, and the diodes were completely blocked for this measurement and put in a dark box. The bias was 10 V at first, but then reduced to 5 V in an attempt to reduce the excess noise after I read on the datasheet that 10 V was an absolute maximum for some reason. I did not record the difference in noise from 10 V to 5 V, but this is a test I will likely try (though perhaps not up to 10 V anymore).

While 7845 is clearly bad, the others are probably OK for now; they are not acceptable for low-power/high-Z operation, but are likely just fine for our high-power testing since we will expect shot noise levels of >100 pA/rtHz, with SNR with respect to PD noise increasing as \sqrt{P}.

  1206   Fri Feb 20 05:41:04 2015 ZachDailyProgressLaserSimple ISS test

I wanted to do some intensity feedback testing, for two reasons:

  1. Just to get used to using the fiber amplitude modulators
  2. While I wait for the machined parts for the 1064nm M2 ISS testing on the old gyro table, I might as well use this basically perfect setup to do some initial runs with the M2 at 1550nm

So that I can have as much power as possible, I removed the fiber phase modulator and installed the amplitude modulator in its place. To generate the PDH sidebands, I simply drove into the laser bias tee with the 30 MHz oscillator signal and increased the amplitude until I got the same modulation depth as I measured with the modulator. I also had to readjust the demod phase via cable lengths, but after that the cavity locked just as before (and with an identical OLTF---not shown here). I don't claim that this locking technique is as good as using a phase modulator, in light of possible RFAM effects, but it is likely fine for intensity testing.

I also tried to increase the DC drive current of the laser, but it kept stalling after I tried to increase it above ~115 mA (the output power would increase in accordance with the plot on the datasheet, but then would suddenly crash and not return if the current was lowered until the driver output ON/OFF was cycled---not sure what gives here). So, I set it to 100 mA, where it seemed stable. The output of the laser head at this current is ~12 mW, so the max-transmission output of the amplitude modulator is about 6 mW (due to the 50% insertion loss). Adding a slight DC offset to the modulator, I reduced the output to ~92% to get some linear actuation strength for feedback.

I then tried to create an AC-coupled loop with an SR560, but had problems with stability on the low end. Eventually, I gave up and used the A-B function to subtract the measured DC level of around 4 V from the TRANS PD signal. I then put a pole at 300 Hz and scaled up the gain until I saw oscillations up near 100 kHz, and then slightly back down. Using this offset-subtracted DC-coupled loop, I was able to get solid in-loop performance, obtaining a UGF near 100 kHz and suppressing fluctuations to the dark noise level (consistent with the PDA255's noise) over a wide band.

The next step will be to use my low-noise readout optoelectronics and try out the Chachi servo.

 

  1205   Fri Feb 20 05:22:19 2015 ZachDailyProgressLaserError calibration -> actuation TFs and new laser frequency noise measurement

Using the REFL PDH setup I built the other day (and that was detailed somewhat by Nic and Chris W. in CRYO:1204), I calibrated the error response so that I could make some further measurements. To refresh, this is using 30-MHz sidebands applied using one of our fiber phase modulators, sensing with a 1611 in reflection. The sideband drive was 0 dBm.

Using the sidebands as a reference, I calculated the slope at 6.4 nV/Hz:

Note that the error signal is slightly asymmetric, but there is no large offset.

With this information, I made some measurements:

Actuation transfer functions

I wanted to measure the actuation transfer functions afforded by:

  1. The standalone ThorLabs laser diode driver (LDC201C) that is currently used to drive the west laser
  2. The diode driver in the integrated ThorLabs laser controller (ITC502) that is currently used to drive the east laser
  3. Simply driving into the bias tee on the diode

All measurements were done with the west laser head, driving as described and reading out at the error point, then correcting for the loop gain and calibrating to Hz.

Everything is more or less as expected:

  • The LDC201C only claims 3 kHz bandwidth, which is actually a bit of a stretch as usual
  • The ITC502 claims 500 kHz, and this is also a stretch (there is -45° at 40 kHz), but it's not so bad for some early locking
  • Directly driving into the diode works across this measurement band and likely far, far beyond
  • There are some common features, prominently at low frequencies and somewhat so at higher ones that likely arise from the current -> frequency response of the diode itself

 

Laser frequency noise

I now have another calibrated measure of laser frequency noise, wherever it dominates over the PMC length noise. I measured the error signal, corrected for the loop gain and calibrated to Hz. For comparison, I've added the measurement using the Zurich PLL on the beat between the two free-running lasers on 12/17/2014 (see CRYO:1185), as well as the RIO spec for this laser.

As you can see, tonight's measurement agrees quite well with the earlier one upt to ~1 kHz, above which the old measurement is probably marred by the relatively low-bandwidth PLL. It seems that the PMC is quiet enough to see the laser noise throughout, and the new measurement now sits closer to the spec up to the highest available point at 10 kHz. Below ~50 Hz, we are probably seeing the well-documented excess noise from the ThorLabs driver. Everything looks as expected.

Locking via laser feedback

Relatively early in the night, after having measured the actuation transfer functions, I sucessfully locked the cavity via feedback to the laser (as opposed to the PMC PZT) for the first time. Below is a comparison of the OLTFs for a 1-kHz loop using the same servo shape (a pole at 1 Hz) using both actuation schemes.

Because of what I had hooked up at the time, I only did this with the (low-bandwidth) LDC201C, so, while the absence of a ~10-kHz resonance is clear, the phase margin is not improved at all (worsened, actually). I only report this as a milestone, and the margin afforded by the ITC502 or by directly driving via the bias tee should be far better.

  1204   Thu Feb 12 15:34:12 2015 Nic, ChrisLaserTransfer Functionshigher bandwidth frequency readout

In order to better measure the effect of this nonlinear current to frequency modulation, we'll need to do Zach's measurement but with much higher drive frequencies. (His measurement was 1kHz).

We'd like to do a full TF of the nonlinear current amplitude modulation path to the laser frequency. There are two effects in Zach's setup that limit the bandwidth of the measurement.

First, is the modulation input of the Marconi, which only reaches 30kHz. We plan to use a mixer to do higher frequency AM of the RF carrier.

The second is the frequency readout. We potentially could PLL the two lasers together and have a pretty high bandwidth readout. or, instead we decided to add some additional PDH sidebands to the light using the fiber modulator. This was then sensed in reflection of the PMC and demodulated. We used 30MHz at 0.5Vpp into the fiber modulator.

With this setup, we were able to measure some amount of nonlinear current to frequency modulation, and when we unlocked the cavity the transfer function was reduced by at least 20dB, which rules out some other coupling path.

Next step is to set up high bandwidth AM of the 500MHz marconi output (driving the current).

  1203   Wed Feb 11 03:58:52 2015 ZachLaserTransfer FunctionsWipf nonlinear temperature actuation proof of principle

Nic elucidated to me today Chris W.'s idea for getting truly wideband (~500 MHz) actuation out of our diode lasers. In case the reader isn't familiar, the lasers have two parallel linear actuation pathways converting current into frequency: one from current modulating the temperature, which is the strongest effect at DC and then dies off above ~1 MHz due most likely to the thermal response, and another, weaker but much wider-band, flat pathway arising from solid state effects that did not survive the elucidating. At some frequency (around 50 MHz, I believe?), there is a crossover between these paths, but there is a differing sign, which creates a "non-minimal-phase zero", leaving the phase at -180° and making the overall system a difficult actuator to deal with at high frequencies.

As I understand it, Chris's idea involves using the full, nonlinear current-to-temperature response to effectively circumvent the direct linear response at low frequencies. This can be done, for example, by pumping a strong RF carrier current (say, around 1 GHz) into the diode, and then using amplitude modulation on this carrier to produce baseband frequency actuation from the temperature beating. By choosing the phase of the AM correctly, one can make it so this pathway (now dominant at low frequencies) results in a nicer crossover with linear pathway #2 from above.

I performed a very simple proof-of-principle test today by doing the following:

  • Dither lock my temporary diagnostic PMC to one laser using the setup described in CRYO:1195.
  • Set the UGF fairly low (a few 100 Hz)
  • Drive the laser current with a 1-kHz sine wave, strong enough to be clearly present above the noise in the error signal. I found that 200 uVpp (= 2 uApp) gave me a nice SNR around 20.
  • Using a Marconi into the SMA bias tee adapter directly on the diode, inject a fairly strong RF carrier current. I used 600 MHz at ~200 uArms, though the amplitude was determined empirically over the course of the test to see an effect.
  • Engage amplitude modulation at 1 kHz and a pretty strong modulation (I chose "50%").
  • (As I mentioned a couple bullets above, in reality, I removed the direct 1-kHz injection and pumped this RF-with-AM current up until I saw an effect in the error signal)
  • With these two signals on, and adjusting the AM phase, I was clearly able to see modulation of the line in the error signal, indicating that the two drives were interfering as desired.

Trimming the RF amplitude and phase a bit to get a nice result, I was able to take the two spectra shown below. In the first trace, only the direct current line is present at 1 kHz. In the second one, the RF source is engaged and you can see an exact cancellation of the line in the error signal. Increasing or decreasing the RF (or audio) amplitudes led to the reemergence of the line (assuredly with 180º relative phase from one case to the other). To do the wideband actuation, one would simply make sure that the RF power is strong enough that the nonlinear path dominates.

So, it should work! We'll have to change the measurement setup to make a full transfer function showing clean actuation to very high frequencies, but it should be pretty straightforward.

  1202   Wed Feb 11 03:15:02 2015 ZachDailyProgressSiFi - ringdownOnly some extra damping is from wires

Following my preliminary conclusion from yesterday (CRYO:1201), I set out to confirm or deny this seeming decrease in Q for a given clamp when going from the simple vacuum chamber to the cryostat.

One potential source of extra damping I considered was the wires attached to the block for the power resistor and RTD, so, while I still had the clamp in the cryostat assembly, I just disconnected these wires and pumped down the cryostat to see if I saw an improvement. I did see an increase in Q from ~3000 to ~5500, but not to the full 7000 I saw before in the standalone chamber. So, I conclude that there is some appreciable damping added by this kapton wiring. We need to use less rigid wire for the last stretch between the coldplate-mounted strain releif and the block.

The last step was to transport the clamp back into the simple chamber and see if I could recover the Q of 7000 that I measured initially. I did, completing the circle of repeatablility. I'm not sure what else could be causing the excess damping in the cryostat.

It is a shame, because I would be very interested to see what this particular silicon sandwich clamp looks like at 120 K, but I seem to have now way of doing so without the extra losses empirically associated with putting it in the cryostat.

  1201   Tue Feb 10 04:37:05 2015 ZachDailyProgressSiFi - ringdownQ consistently lower in cryostat

A lot of things happened tonight (mostly in the realm of setbacks followed by recovering frome them), but the take-home is that the measured Q of my silicon sandwich clamp seems consistently lower when measured in the cryostat, compared to in the new chamber from the gyro. Here's a rundown of what happened today/tonight:

  • Before dinner, I made a first measurement on the silicon sandwich idea (cantilever sandwiched between a couple spare pieces of silicon on each side --- see CRYO:1200). This gave me the highest room-temperature Q I've measured yet at ~6800.
  • After dinner, I wanted to port this to the cryostat and potentially do a cooling run. Unfortunately, to fit it in the cryo volume, I had to flip the sandwich around so that it was protruding from the clamp in the other direction (for the first run, I had it sticking out over the power resistor to avoid clamping in the region on the other side that has the groove for the Glasgow-style cantilevers, but there wasn't enough room for that orientation in the cryostat, so I had to flip back---I made it work so I didn't clamp over the groove anyhow).
  • Unwittingly, I made the dumb mistake of not first testing this freshly-clamped system again in the simple chamber, and after I closed the whole cryostat again and pumped down, I measured a much lower Q (back down around 3000).
  • So, I opened the cryostat again, and then spaced out and made the further mistake of still not testing this apparently bad clamp job in the simple chamber, just to verify that I got the same low Q. Instead, I went straight to cleaning all the pieces and re-clamping.
  • This time, I put it into the simple chamber and immediately recorded a high Q around 7000 again.
  • This is when some setbacks kicked in:
    • In opening the chamber, one of the RTD wires came loose from the feedthrough.
    • Not realizing that these were just press-fit sockets, I unscrewed the feedthrough to have access so I could reattach the single loose wire, only to have several others fall off.
    • So, I disconnected all the wires, spent some time mapping which one went where, re-soldered some and re-kapton shieled all, then reattached all wires, bunch taped them and taped the bunch to the feedthrough so that none could easily come loose. I also took this time to resolder the ESD wire that I broke the other day.
    • In moving stuff around, I accidentally tugged on the ribbon cable between the QPD and its vectorboard readout circuit, pulling a couple connections.
    • So I spent some time fixing that
  • Now I was ready to do science again, so I transferred the (known good) clamp from the simple chamber back into the cryostat and carefully closed it all up again.
  • After seal and pumpdown, I again measured a low Q around 3000.

So, it seems that the Q is repeatably lower for a particular clamp in the cryostat vs. in the simple chamber. To be sure, I'm going to do the final step of returning the clamp back to the simple chamber tomorrow and see if I again get a higher Q.

I'm not exactly sure why this could be happening. The only mechanical differences from one chamber to the other are:

  1. The clamping block is screwed via holes in the PEEK base to the cold plate in the cryostat, while it is dogclamped to the breadboard in the simple chamber.
  2. In the cryostat, there are wires soldered to the power resistor attached to the clamping block as well as a wire-attached Pt RTD kapton-taped to it. None of this is present in the simple chamber.

I'm tempted to think that (2) could be causing some excess damping, so one thing I will try is simply not connecting these just to see if that makes the probem go away.

  1200   Mon Feb 9 18:49:55 2015 ZachDailyProgressSiFi - ringdownQ ~ 6800 at room temp with Si sandwich

As I planned yesterday (CRYO:1199), I tried out a new clamp using spare pieces of broken silicon instead of sapphire washers to sandwich the cantilever (as with the last run, I used the old, stiff rectangular block clamp---the newer cylindrical one is still in the cryostat).

I didn't take a photo, but this was basically just a sandwich consisting of the cantilever (still attached to the central wafer region) as the meat and two scrap broken-off cantilevers on each side as the bread. This was all put near the center of the steel block clamp so that the clamping force was normal, and I made sure that the protruding cantilever had enough room not to be clipped by the block as it swings.

I put it in the new chamber and pumped down, and immediately measured a fairly high Q of ~6800 (ringdown tau ~ 6.4 s, while the mode frequency is ~340 Hz---slightly higher than before due to the clamping being a bit further along the cantilever).

This is the highest room-temperature Q I've yet measured, beating the ~4300 I measured after we first installed the sapphire washers on the newer cylindrical clamp (see CRYO:1191), and is within a factor of 2 of Marie's prediction in the absence of clamping loss (also shown in that post). This is also by far the cleanest ringdown I've seen: there are a few high-frequency modes present when I first deliver the impulse, but they die away and do not return. The Q also seems far less amplitude-dependent than I've noticed before.

ELOG V3.1.3-